A hiker began a hike with 0.86 gal water. At the end of the hike, 721.8 mL of water remained. What percent of the water was used during the hike? Type your numerical answer to 1 decimal place (ex NN.N) without the % sign.
1
Expert's answer
2017-09-04T14:49:06-0400
1 gal = 3 785 ml
0.86 gal = 0.86 × 3785 ml = 3255 ml
ΔV = 3255 ml – 721.8 ml = 2533.2 ml
%v/v=(2533.2 ml)/(3255 ml)×100%=77.8%
Answer: 77.8%.